You are considering an investment in Justus Corporation's stock, which is expected to pay a dividend of $2.00 a share at the end of the year (D1 = $2.00) and has a beta of 0.9. The risk-free rate is 5.8%, and the market risk premium is 5.0%. Justus currently sells for $21.00 a share, and its dividend is expected to grow at some constant rate, g. Assuming the market is in equilibrium, what does the market believe will be the stock price at the end of 3 years? (That is, what is ?) Round your answer to two decimal places. Do not round your intermediate calculations.

$  
Answer #1

Cost of equity using CAPM method =Risk Free Rate+Beta*Market Risk Premium =5.8%+0.9*5% =10.30%
Current Share Price =Dividend Year 1/(Required Rate-Growth) =2*(1+g)/(10.30%-g) =21
2+2g =21*10.30%-21g
g =(21*10.30%-2)/23 =0.7089%

Price in year 3 =P0*(1+g)^3 =21*(1+0.7089%)^3 =21.45